what mass in grams are in 5.32*10^22 molecules of co2


1
Expert's answer
2022-03-08T14:05:03-0500

1 mol = 6.022*1023 number of molecules.

n(CO2) = 5.32*1022/6.022*1023 = 0.0883 mpl

m(CO2) = 0.0883 * 44 = 3.89 g

Answer: 5.32*1022 molecules of CO2 is 3.89 g